OpenStudy (kayleahjb):

ook at the right triangle ABC: Right triangle ABC has a right angle at B. Segment BD meets segment AC at a right angle. A student made the following chart to prove that AB2 + BC2 = AC2: Which justification is incorrect? (1 point) Justification 4 Justification 1 Justification 2 Justification 3

OpenStudy (kayleahjb):

Statement Justification 1. Triangle ABC is similar to triangle BDC 1. Angle ABC = Angle BDC and Angle BCA = Angle BCD 2. BC2 = AC ⋅ DC 2. BC ÷ DC = AC ÷ BC because triangle ABC is similar to triangle BDC 3. Triangle ABC is similar to triangle ABD 3. Angle ABC = Angle BAD and Angle BAC = Angle ABD 4. AB2 = AC ⋅ AD 4. AB ÷ AD = AC ÷ AB because triangle ABC is similar to triangle ADB 5. AB2 + BC2 = AC ⋅ AD + AC ⋅ DC = AC (AD + DC) 5. Adding Statement 1 and Statement 2 6. AB2 + BC2 = AC2 6. AD + DC = AC

OpenStudy (kayleahjb):

I think D

OpenStudy (error1603):

its A